OpenStudy (anonymous):

find the derivative of y=log10(5x^4-2/x^3)

OpenStudy (anonymous):

this involves both the chain rule and the quotient rule

OpenStudy (mimi_x3):

or you can use log laws to change it first; then diferentiate it, would be easier.

OpenStudy (anonymous):

@failmathmajor only the 2/x^3 is a fraction

OpenStudy (mimi_x3):

is it \[\log\left(5x^{4}-\frac{2}{x^{3}}\right) ?\]

OpenStudy (anonymous):

that would be much easier

OpenStudy (anonymous):

so using log properties you can say that \[\log_{10} {x} = \ln x / \ln 10 \]

OpenStudy (anonymous):

ln10 is a constant, you can factor that out

OpenStudy (anonymous):

can you use the chain rule fluently?

OpenStudy (anonymous):

y'=pu(x) * u'(x) right?

OpenStudy (anonymous):

d(f(g(x))/dx = (dg(x)/dx)*(df(g(x))/dx) is the way I know it

OpenStudy (mimi_x3):

Or..\[y= \log\left(5x^{4}-\frac{2}{x^{3}}\right) y = \log\left(\frac{5x^{7}-2}{x^{3}}\right) => y = \log\left(5x^{7}-2\right) - \log(x^3)\] Might be easier; than the quotient role and chain rule..

OpenStudy (anonymous):

that may look confusing now that I see it

OpenStudy (anonymous):

don't need the quotient in this case: can treat /x^3 as a x^-3

OpenStudy (mimi_x3):

@cunninnc: are you able to do it now?

OpenStudy (anonymous):

and you would still need the chain rule for that anyway

OpenStudy (anonymous):

but enough arguing

OpenStudy (mimi_x3):

Did an argument started? lol

OpenStudy (anonymous):

\[(d(\ln (5x^4 - 2x^{-3}))/dx )/ \ln10\] is what it simplifies down to, to be concise

OpenStudy (anonymous):

@Mimi_x3 kinda .... i see mr. moose got 2x^-3 where does ^-3 come froms

OpenStudy (anonymous):

\[2/x^3= 2 * x^{-3}\]

OpenStudy (mimi_x3):

Sorry i don't know what MrMoose is doing. @MrMoose: Use \frac{x}{y} for fractions :)

OpenStudy (mimi_x3):

Or why not try the method that i used :) \[ \frac{d}{dx} \log\left(5x^{7}-2\right) -\frac{d}{dx} \log(x^3)\]

OpenStudy (anonymous):

when you divide you subtract exponents, so that is equivalent to saying \[2 * \frac{x^0}{x^3}\] then subtract exponents in division

OpenStudy (anonymous):

\[\frac{\frac{d(\ln(5x^4−2x^{−3}))}{dx}}{\ln10}\]

OpenStudy (anonymous):

@catamountz15 what?

OpenStudy (anonymous):

I am almost entirely sure that that isn't a form of the answer.

OpenStudy (anonymous):

Here are the steps in to solving this problem.

OpenStudy (anonymous):

that isn't what you wrote though

OpenStudy (anonymous):

My apologies that was an answer to a different problem.